2.) Acetone is used in a lot industrial processes. It is highly flammable and easily dissolves in water. A mixture of acetone and water is made of 75% water and 25% acetone. The solution is run through a chiller at -100 degrees Celsius. At this temperature, water will freeze but not acetone. The mixture is fed in at 350 gal/min, and 125 gal/min of acetone is added to maintain temperature. How much acetone is recovered from the chiller?
Answer #1

Ans.

given,

the mixture is fed in at 350 gal/min.

and given that the mixture of acetone and water is made of 75% water and 25% acetone.

hence, acetone fed by mixture is given by 25% of mixture.

25 a1 - 350 X 100 = 87.5 gal/min

and given that additional acetone added to mantain teperature is

\large a_{2}=125 \ gal/min.

hence, total acetone recovered from chiller is

\large a=a_{1}+a_{2}=125+87.5=212.5 \ gal/min.